Hello, Could you please help me figure out if I have the problem
8ft. 4in.-2ft.9in. I have to borrow from the 8 but I can't figure out how many inches add on to the 4 to be able to subtracted from the 9. Thanks

Answer by ptaylor(2198)

ONE APPROACH WOULD BE TO CONVERT EVERYTHING TO INCHES AND THEN CONVERT BACK:
8FT 4IN =96+4=100 IN
2FT 9IN =24+9=33IN
100 IN -33 IN=67 IN
67/12 =5 FT 7 IN
Another way
8ft 4 in=7ft 16 in minus 2 ft 9 in =5ft 7 in
Hope this helps---ptaylor